one positive number exceeds twice another number by 3 and the product of the two numbers is 44. Find the smaller number.
Answer by CubeyThePenguin(3113)

x = smaller number
y = larger number

y = 2x + 3
xy = 44

Substitute.
x(2x + 3) = 44
2x^2 + 3x - 44 = 0
(2x + 11)(x - 4) = 0

The numbers are positive, so smaller = 4 and larger = 11.